Check the function f(x)=[2x-1, 0≤x≤1
[1, x>1
for continuity and differentiability at x =1
1
Expert's answer
2019-04-16T09:43:34-0400

"\lim_{x\to 1^-}f(x)=\lim_{x\to 1^+}f(x)=1\:\;so,\;f(x)\;is\;continuous\;at\;x=1\\
\lim_{\to 1^-}\frac{f(x)-f(1)}{x-1}=\lim_{\to 1^-}\frac{2(x-1)}{x-1}=2,
\lim_{\to 1^+}\frac{f(x)-f(1)}{x-1}=0\\
so,\;f(x)\;is\;not\;differentiable\;at\;x=1."